OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (-1,-7)(3,5)

OpenStudy (anonymous):

1st of all, do u know the formula of line equation ? tell me

OpenStudy (anonymous):

(y-y1)=((y2-y1)/(x2-x1))(x-x1) y+7=(5+7/3+1)(x+1) y+7=(12/4)(x+1) y+7=3(x+1) y+7=3x+3 3x-y=4